A company is projected to generate free cash flows of $443 million next year, growing at a 4.6% rate until the end of year 3. After that, cash flows are expected to grow at a stable rate of 2.7%. The company's cost of capital is 10.4%. The company owes $257 million to lenders and has $26 million in cash. If it has 171 million shares outstanding, what is your estimate for its stock price? Round to one decimal place.
Answer #1

Expected FCF, FCF1 = $443,000,000

Growth rate for next 2 years is 4.60% and a constant growth rate of 2.70% thereafter

FCF2 = $443,000,000 * 1.046 = $463,378,000
FCF3 = $463,378,000 * 1.046 = $484,693,388
FCF4 = $484,693,388 * 1.027 = $497,780,109.476

Cost of Capital = 10.40%

Horizon Value of Firm at the end of Year 3 = FCF4 / (Cost of Capital - Growth Rate)
Horizon Value of Firm at the end of Year 3 = $497,780,109.476 / (0.1040 - 0.0270)
Horizon Value of Firm at the end of Year 3 = $6,464,676,746.44

Value of Firm = $443,000,000/1.104 + $463,378,000/1.104^2 + $484,693,388/1.104^3 + $6,464,676,746.44/1.104^3
Value of Firm = $5,946,073,340.58

Value of Equity = Value of Firm - Value of Debt + Value of Cash
Value of Equity = $5,946,073,340.58 - $257,000,000 + $26,000,000
Value of Equity = $5,715,073,340.58

Stock Price per share = Value of Equity / Number of Shares Outstanding
Stock Price per share = $5,715,073,340.58 / 171,000,000
Stock Price per share = $33.42